# Structured Probabilistic Coding

### v4

@@ Introduction
We conduct experiments on 12 natural language understanding tasks, including 10 classification tasks such as emoji prediction, hate speech detection, irony detection, offensive language detection, sentiment analysis, stance detection, emotion detection from different domains, as well as 2 regression tasks including semantic similarity prediction and plausible clarifications ranking. The results demonstrate that our SPC effectively improves the performance of pre-trained language models for classification and regression tasks. For instance, with the RoBERTa backbone, SPC improves average performance by +4.0% and +1.7% for classification and regression tasks compared to CE/MSE. Our SPC framework consistently achieves the best average performance compared to other methods, including deterministic and probabilistic approaches, under different backbone models such as BERT and RoBERTa. Extensive experiments show that SPC can enhance the model generalization capability including out-of-distribution and data-constrained scenarios, robustness to label noise, and clustering quality of output representations.
@@
All experiments are conducted on a single NVIDIA Tesla A100 80GB card. The validation sets are used to tune hyperparameters and choose the optimal model. For each method, we run five random seeds and report the average result of the test sets. The network parameters are optimized by using Adamax optimizer (Kingma and Ba 2015) with the learning rate of $5e^{-5}$, the weight decay coefficient of $\{0,0.01,0.001\}$. The dropout rate is set to 0.2. The trade-off parameter $\beta$ is searched from $\{0.001,0.01,0.1,1,10\}$, and the parameter $\gamma$ is searched from $\{0.01,0.05,0.1,0.5,1,5,10\}$. More experimental details are listed in the Appendix.
@@
SPC $*{}^{*}$$\pm$0.48 79.01$\pm$0.61 $*{}^{*}$$\pm$1.32 65.31$\pm$1.91 80.98$\pm$1.36 72.96$\pm$0.22 $*{}^{*}$$\pm$0.63 $*{}^{*}$$\pm$0.59 $*{}^{*}$$\pm$0.29 $*{}^{*}$$\pm$0.60 62.47
Table 1: Classification evaluation (%) on 10 benchmark datasets. ${}^{\dagger}$ means the results are from Barbieri et al. (2020). For other methods, we run five random seeds and report the average result on test sets. BERT and RoBERTa are the backbone models for deep representation learning technologies. Best results for each dataset are highlighted in bold. $*{}^{*}$ represents statistical significance over state-of-the-art scores under the $t$ test ($p<0.05$).
@@
The overall results for 10 classification tasks are summarized in Table 1. Our SPC consistently obtains the best average performance over comparison methods. When using BERT and RoBERTa backbones, SPC can enhance average performance by +3.1% and +4.0% compared to CE for all classification tasks, respectively. The results indicate the good generalization ability of our method to unseen test sets and show the superiority on classification tasks. We notice that SPC achieves big improvements for HatEval and IronyEval, i.e., +14.3% macro-F1 scores and +7.3% F1 scores of the ironic class, respectively. In HateEval, there is a topic distribution disparity between the validation set and the test set. Additionally, the IronyEval task requires complex semantic understanding, with subtle differences between ironic and non-ironic texts. These results indicate that our SPC has a good generalization capability on the above specific scenarios, i.e., topic shifts and subtle semantic labels.
@@
We experiment to demonstrate the robustness by assessing how well the models can handle noisy labels. It is crucial for real-world scenarios where data can often be unreliable. Specifically, we randomly choose 10%, 20%, 30% of training data and flip their labels to any category randomly with equal probability. We run experiments five times and compute the mean and standard variance of the results. As shown in Table 5, under all settings, SPC consistently outperforms CE, VIB, and MEIB. It indicates that SPC performs more robustly on noisy training data. Besides, compared to CE, SPC improves average performance by +2.0%, +2.1%, and +1.7% with noise ratio of 10%, 20%, and 30% on classification tasks. The results prove that SPC can better control and utilize randomness and uncertainty of data.
@@ Appendix A Experimental Setups
Trade-off weight $\beta$ 0.1 0.1 10 1 1 0.01 1 0.1 1 0.001
Trade-off weight $\gamma$ 10 0.1 0.1 10 0.1 10 0.5 0.1 0.05 0.01
Weight decay 0 0 0.001 0 0 0 0 0.001 0 0
Layer normalization False True False True False False False False True False
@@
We conduct experiments using an epoch number of $20$, a total batch size of $128$, and a maximum token length of $128$. The maximum patience for early stopping is set to 5 epochs. We report the detailed hyperparameter settings of SPC with RoBERTa backbone in Table 7 and Table 8. In the implementation of SPC, the hidden vector represents the output representation, and its dimension size is consistent with the dimension size of label space in each task.
@@ Appendix B Supplementary Experiments
In this part, we analyze what the trade-off parameters $\beta$ and $\gamma$ control in our SPC.
@@
Figure 5 shows relative improvements between SPC and its ablated variant (i.e., w/o Structured) against the optimal $\gamma$. By introducing task-related underlying structured patterns in representation learning ($\gamma>0$), SPC achieves varying degrees of relative improvements on all tasks, particularly in HatEval and IronyEval. A larger value of $\gamma$ indicates that this type of task requires enhanced task-related learning ability.

### v5

@@ Introduction
We conduct experiments on 12 natural language understanding tasks, including 10 classification tasks such as emoji prediction, hate speech detection, irony detection, offensive language detection, sentiment analysis, stance detection, emotion detection from different domains, as well as 2 regression tasks including semantic similarity prediction and plausible clarifications ranking. The results demonstrate that our SPC effectively improves the performance of pre-trained language models for classification and regression tasks. For instance, with the RoBERTa backbone, SPC improves average performance by +3.9% and +1.7% for classification and regression tasks compared to CE/MSE. Our SPC framework consistently achieves the best average performance compared to other methods, including deterministic and probabilistic approaches, under different backbone models such as BERT and RoBERTa. Extensive experiments show that SPC can enhance the model generalization capability including out-of-distribution and data-constrained scenarios, robustness to label noise, and clustering quality of output representations.
@@
All experiments are conducted on a single NVIDIA Tesla A100 80GB card. The validation sets are used to tune hyperparameters and choose the optimal model. For each method, we run five random seeds and report the average result of the test sets. Besides, we conduct experiments using an epoch number of $20$, a total batch size of $128$, and a maximum token length of $128$. The maximum patience for early stopping is set to 5 epochs. The network parameters are optimized by using Adamax optimizer (Kingma and Ba 2015) with the learning rate of $5e^{-5}$, the weight decay coefficient of $\{0,0.01,0.001\}$. For SPC, the trade-off parameter $\beta$ and $\gamma$ are searched from $\{0.001,0.01,0.1,1,10\}$ respectively. More experimental details are listed in the Appendix.
@@
SPC 32.54$\pm$0.48 79.01$\pm$0.61 59.80$\pm$1.32 65.31$\pm$1.91 80.98$\pm$1.36 72.96$\pm$0.22 69.02$\pm$0.63 71.01$\pm$0.59 43.99$\pm$0.29 48.92$\pm$1.83 62.35
Table 1: Classification evaluation (%) on 10 benchmark datasets. means the results are from Barbieri et al. (2020). For other methods, we run five random seeds and report the average result on test sets. BERT and RoBERTa are the backbone models for deep representation learning technologies. Best results for each dataset are highlighted in bold. represents statistical significance over state-of-the-art scores under the $t$ test ($p<0.05$).
@@
The overall results for 10 classification tasks are summarized in Table 1. Our SPC consistently obtains the best average performance over comparison methods. When using BERT and RoBERTa backbones, SPC can enhance average performance by +3.1% and +3.9% compared to CE for all classification tasks, respectively. The results indicate the good generalization ability of our method to unseen test sets and show the superiority on classification tasks. We notice that SPC achieves big improvements for HatEval and IronyEval, i.e., +14.3% macro-F1 scores and +7.3% F1 scores of the ironic class, respectively. In HateEval, there is a topic distribution disparity between the validation set and the test set. Additionally, the IronyEval task requires complex semantic understanding, with subtle differences between ironic and non-ironic texts. These results indicate that our SPC has a good generalization capability on the above specific scenarios, i.e., topic shifts and subtle semantic labels.
@@
We experiment to demonstrate the robustness by assessing how well the models can handle noisy labels. It is crucial for real-world scenarios where data can often be unreliable. Specifically, we randomly choose 10%, 20%, 30% of training data and flip their labels to any category randomly with equal probability. We run experiments five times and compute the mean and standard variance of the results. As shown in Table 5, under all settings, SPC consistently outperforms CE, VIB, and MEIB. It indicates that SPC performs more robustly on noisy training data. Besides, compared to CE, SPC improves average performance by +1.9%, +2.2%, and +1.6% with noise ratio of 10%, 20%, and 30% on classification tasks. The results prove that SPC can better control and utilize randomness and uncertainty of data.
@@ Appendix A Experimental Setups
Trade-off weight $\beta$ 1 0.1 10 1 0.01 0.01 1 0.1 1 0.001
Trade-off weight $\gamma$ 10 1 0.1 0.1 0.1 10 0.1 0.1 0.1 0.001
Weight decay 0 0 0.01 0 0 0 0.001 0.001 0 0
Dropout 0 0 0 0.2 0.2 0 0.2 0 0 0.2
Layer normalization True True False True True False True False True False
@@
We report the detailed hyperparameter settings of SPC with RoBERTa backbone in Table 7 and Table 8. In the implementation of SPC, the hidden vector represents the output representation, and its dimension size is consistent with the dimension size of label space in each task.
@@ Appendix B Supplementary Experiments
In this part, we analyze what the trade-off weights $\beta$ and $\gamma$ control in our SPC.
@@
Figure 5 shows relative improvements between SPC and its ablated variant (i.e., w/o Structured) against the optimal $\gamma$. By introducing the underlying structured patterns related to the target task ($\gamma>0$), SPC achieves varying degrees of relative improvements on all tasks, particularly in HatEval and IronyEval. A larger value of $\gamma$ indicates that this type of task requires enhanced task-related learning ability.